# Rhapsody: A Dataset for Highlight Detection in Podcasts

### v1

@@ Abstract
Podcasts have become daily companions for half a billion users. Given the enormous amount of podcast content available, highlights provide a valuable signal that helps viewers get the gist of an episode and decide if they want to invest in listening to it in its entirety. However, identifying highlights automatically is challenging due to the unstructured and long-form nature of the content. We introduce Rhapsody, a dataset of 13K podcast episodes paired with segment-level highlight scores derived from YouTube’s ‘most replayed’ feature. We frame the podcast highlight detection as a segment-level binary classification task. We explore various baseline approaches, including zero-shot prompting of language models and lightweight finetuned language models using segment-level classification heads. Our experimental results indicate that even state-of-the-art language models like GPT-4o and Gemini struggle with this task, while models finetuned with in-domain data significantly outperform their zero-shot performance. The finetuned model benefits from leveraging both speech signal features and transcripts. These findings highlight the challenges for fine-grained information access in long-form spoken media. We release our codes and dataset at https://github.com/younghanstark/rhapsody.
@@ 1 Introduction
Podcast consumption has grown significantly and recent work has studied podcast content-related tasks such as search (Tian et al., 2022), recommendation (Fan et al., 2023; Nadai et al., 2024), and summarization (Jones et al., 2020; Karlgren et al., 2021). In this paper, we present the first study on locating the most interesting parts (i.e., highlights) of podcast episodes automatically, which can be useful for listeners. Because podcast episodes are often lengthy, it is not easy to quickly locate such highlights. Furthermore, what listeners may consider to be the most interesting parts of the podcast is ill-defined and subjective in nature, making it challenging to gather highlight annotations.
Recently, YouTube introduced the ‘most replayed’ feature: a graph displaying frequently replayed parts of videos (referred to as the replay graph) (YouTube, 2022). This feature, which aggregates the views of thousands of users, opened doors for researchers to frame and study video understanding tasks (Duico et al., 2023; Sul et al., 2023; Kim et al., 2025). Inspired by these works, we study podcasts using this replay graph feature. Compared to existing works which mainly capture visual features, our work focuses on the spoken content.
@@
We derive Rhapsody, [^1] Rhapsody: Replay graph, Highlights, Podcasts, Segment-level, Detection. a large-scale dataset (containing 13K podcast episodes from various domains) and propose the podcast highlight detection task, visualized in Figure 1. This task is long-context in nature because the average podcast episode in our dataset is about $30$ minutes long and contains an average of 5K words. Each podcast is divided into $100$ equal-length segments, of which a subset (with an average of $5$ segments) is identified as being the most-replayed highlights. We use YouTube’s replay graph feature to identify these most-replayed segments. Given input features derived from the podcast episode, such as episode transcript, episode title, or audio features, a model must identify this subset of most-replayed segments.
To tackle this new task, we propose a suite of baseline systems. We first present a text-based BM25 (Robertson & Zaragoza, 2009) approach, which treats highlight detection as the task of retrieving the segments whose transcript summaries (as summarized by an off-the-shelf LLM, GPT-4o mini (OpenAI, 2024b)) have the highest BM25 score to the episode title. We next present zero-shot prompting results for two state-of-the-art large language models (OpenAI, 2024a; Google Deepmind, 2024). Lastly, we present learned models on this task, which uses hidden representation from a finetuned LLM (Meta, 2024) and optionally audio embeddings. This model outperforms GPT-4o’s and Gemini’s zero-shot performance by a large margin with only $\sim 6.7$M trainable parameters. The task remains very challenging, with a maximum hit rate of $47.7\%$ with our best approach.
@@
  • We formally describe the podcast highlight detection task, explain how supervision signals can be generated from raw data, and propose a set of automatic evaluation metrics for this task.
  • We benchmark the performance of feature-agnostic and retrieval-based methods, zero-shot LLM prompting, and our finetuned language model.
We release all our data and code to support future research on podcast understanding. Given the limitation of current models, our benchmark can support rich future work, including methods that combine text and audio signals and methods that focus on the long-context ability of LLMs.
@@ 2.1 Task Formulation
We propose a podcast highlight detection task where a model is provided input features derived from the podcast episode and is asked to predict a set of highlight indices $P=\{p_{k}\}_{k=1}^{N_{P}}$ (where $p_{k}\in\{1,2,\dots,100\}$) which is compared to the GT highlight segment indices $G$. We consider both episode-level input features, such as the episode title, and segment-level input features, such as transcript summaries or audio embeddings for each segment.
@@ 2.2 Data Collection Process
We scrape the list of most popular podcast creators from YouTube, [^2] https://www.youtube.com/podcasts/popularcreators yielding 100 channel IDs. Using YouTube Data API v3, [^3] https://developers.google.com/youtube/v3 we then collect metadata for these channels and their playlists. We discard playlists that (1) are not explicitly set as podcasts by their creators and (2) are not in English. We further filter out playlists that predominantly feature visual content, such as music videos, TV series reviews, or reaction videos. We also discard videos for which the replay graph is unavailable. Lastly, we limit the maximum duration of the episode to an hour to prevent a single data point in a replay graph corresponding to a long span of the audio. Details on filtering channels and playlists are in Appendix A.1.
@@
Figure 2: Comparing two settings for defining GT highlights indices $G$. (Left) Naively defining segments with top-$K$ replay scores as highlights. (Right) Introducing $\tau$, hyperparameter to impose a minimum threshold (dotted green lines). The light blue lines at the beginning of the graphs represent the graph before bias correction (Appendix A.3).
@@
3 Comparison Systems
We compare the performance of eight systems: two feature-agnostic methods, a BM25-based method, three zero-shot prompting methods with GPT-4o and Gemini, and two finetuned LLMs—one with a segment-level binary classification head and the other additionally incorporating audio features. We first describe extracting features from raw audio that will be used by systems.
@@ 3.1 Input Data Preprocessing
To map the podcast transcript to the replay graph, which consists of 100 numbers, we split the transcript into 100 segments. We first run the podcast audio through WhisperX (Bain et al., 2023), a version of Whisper (Radford et al., 2022) that outputs a time-aligned ASR transcription providing word-level alignments of the form $\{(w_{j},s_{j},e_{j})\}_{j=1}^{N_{w}}$, where $w_{j}$ represents the $j$th word, $s_{j}$ its start time, $e_{j}$ its end time, and $N_{w}$ the total number of words. Given this WhisperX output, we segment the transcript into 100 equal-duration intervals based on the episode segments $\{c_{i}\}_{i=1}^{100}$. A word $w_{j}$ is assigned to a segment if both $s_{j}$ and $e_{j}$ fall entirely within that segment’s time span. If a word spans two segments, it is assigned to the segment with the greater overlap. After splitting the transcript, we discard episodes with more than 10 segments that do not include any words. We use whisper-large-v2 for the transcription model and WAV2VEC2_ASR_LARGE_LV60K_960H (Baevski et al., 2020) for the alignment model.
@@
We summarize the transcript segments using an LM (gpt-4o-mini-2024-07-18 (OpenAI, 2024b)) and use the summary output as input textual features to the model. We prompt LLM to process 10 consecutive segments at a time while providing the previous context. We provide more detailed information along with pseudocode in Appendix B.1
@@
To capture emotional characteristics in podcasts, we extract dominance-valence-arousal (DVA) embeddings for each podcast segment. The DVA model represents emotions along three dimensions—dominance, valence, and arousal—providing a structured framework for analyzing affective states in speech (Russell & Mehrabian, 1977). When comparing the emotion score distributions of highlight and non-highlight segments, we observe that highlight segments have slightly higher mean scores across all dimensions. To learn richer latent patterns, we leverage the hidden representation from the final transformer layer of an off-the-shelf DVA-based emotion recognition model (Wagner et al., 2023), specifically audeering/wav2vec2-large-robust-12-ft-emotion-msp-dim. A more detailed explanation of audio feature extraction can be found in Appendix B.2.
@@ 3.2 Models
As feature-agnostic baselines, we consider random sampling and frequency-based baseline. The random sampling method samples $K$ unique elements from the population of 100 integers $\{1,2,\cdots,100\}$. The frequency-based method precomputes the occurrence count of each index across all GT highlight indices in the training set and selects the top-$K$ most frequent ones. In both methods, $K$ is sampled from the distribution of $|G|$, the number of GT indices, in the training set. All results from these feature-agnostic baselines were averaged over 1,000 trials.
@@
As a text-only baseline, we adapt BM25 for our task. We treat the summaries of each transcript segment as the corpus and the episode title as the query. Segments with a BM25 score above 3.0 are selected as predicted highlights while preserving the order of the original score. This threshold is determined through manual tuning on the validation set.
@@
To test state-of-the-art LLMs on this long-context reasoning task, we consider $3$ zero-shot settings: GPT-4o (text-only), Gemini (text-only), and Gemini (text + audio). For the GPT-4o (text-only) and Gemini (text-only) settings, we prompt gpt-4o-2024-08-06 and gemini-2.0-flash-001 with the title of the episode and transcript summaries for each segment, asking them to identify the segment indices that are most likely to be replayed. The model is instructed to rank answer segments in descending order of replay likelihood, and chain-of-thought prompting is applied in the prompt. For the Gemini (text + audio) setup, we also provide the full podcast audio. The text-only and the text+audio prompts can be found in Appendix C.2.
Finetuned LLMs with Segment-Level Classification Heads
@@
We finetune an LLM on Rhapsody using parameter-efficient finetuning and replacing the default next-token prediction head with a custom segment-level binary classification head, shown in Figure 3. We develop two models, a text-only model that only utilizes textual features and a text + audio model that utilizes both textual and audio features. The input to the text-only model is $100$ transcript summaries, one for each of the $100$ segments, with end-of-segment tokens appended after each segment. [^4] Specifically, we use one of Llama’s reserved special tokens, <|reserved_special_token_0|>. We extract one LLM hidden representation for each of the $100$ end-of-segment tokens and apply a binary classification head, consisting of a linear layer and a sigmoid activation, to obtain $100$ probability values, each estimating the likelihood that the corresponding segment is a highlight. In the text + audio setup, for each of the $100$ audio segments, we extract one VDA (valence-dominance-arousal) embedding using an off-the-shelf classifier (Wagner et al., 2023), apply a linear transform and concatenate it with its corresponding LLM hidden representation.
We freeze the base model, add QLoRA (Hu et al., 2022; Dettmers et al., 2023) adapters and optimize these adapters as well as the classification head and linear layer of audio features using binary cross-entropy loss. We use Llama-3.2-1B-Instruct as the base model and apply AdamW optimization with weight decay. The best hyperparameter setting (Appendix D) by manual tuning on the validation split results in $\sim$5.6M and $\sim$6.7M trainable parameters in text-only and text and audio settings, respectively.
@@ 4.1 Evaluation Metrics
We aim to measure the similarity of a sequence of predicted indices $P=\{p_{k}\}_{k=1}^{N_{P}}$ based on a sequence of GT indices $G=\{g_{j}\}_{j=1}^{N_{G}}$. Since this setting can be viewed as a document retrieval task where $P$ is a set of retrieved documents and $G$ is a set of relevant documents, we adopt five widely used evaluation metrics for document retrieval — hit, precision, recall, F1, and average precision (AP). When computing AP, $P$ is assumed to be sorted in descending order of confidence to be a highlight according to the model, and we ensure $m_{g_{i}}\geq m_{g_{j}}$ is satisfied for all $1\leq i<j\leq N_{G}$.
The number of predictions heavily influences many evaluation metrics, such as precision and recall, and maintaining a comparable number of predictions is crucial for a fair comparison. Therefore, we tune the hyperparameters on the validation set so that all baseline models produce a similar average number of predictions. All non-baseline models also maintain a different but internally consistent average number of predictions, except for zero-shot prompting methods, which cannot be controlled.
@@ 4.2 Results
System Hit Prec. Recall F1 AP
@@
Finetuned Llama (text-only) 0.452 0.202 0.178 0.163 0.134
Finetuned Llama (text + audio) 0.477 0.216 0.187 0.175 0.140
@@
We report the main experimental results in Table 2. Our experimental results show that even the best-performing model achieves only 0.216 precision and 0.175 F1 score, highlighting the difficulty of podcast highlight detection and the room for improvement beyond current approaches. Yet, this is a subjective task with derived gold labels, thus the upperbound performance might not be perfect. We still anticipate a large room for improvement given the current performance level.
As expected, random sampling and frequency-based approaches performed the worst. Amongst the zero-shot prompting approaches, the Gemini model performed the best, while GPT-4o performed similarly to the frequency-based baseline, suggesting that it struggles to infer highlight segments from textual content without explicit training. Finally, our fine-tuned LLM with a segment-level classification head trained on in-domain data, despite having only $\sim 6.7$M trainable parameters, significantly outperforms all baselines across all metrics. However, the absolute performance remains lacking, emphasizing the need for further advancements in this task.
@@
Incorporating audio features alongside text leads to consistent performance improvements for the finetuned Llama model, suggesting that prosodic and acoustic cues provide valuable information for identifying highlights. Yet, the zero-shot Gemini result with audio does not consistently outperform the text-only Gemini result. This suggests that, while there are useful signals in the spoken content, using it effectively without fine-tuning on in-domain data in a zero-shot manner might be nontrivial. Providing 100 segmented audio clips instead of an entire episode might contribute to the improvement of zero-shot Gemini performance.
@@
FT Llama (text + audio) 3.96 (3.46)
[…]
tableThe average number of GT highlights and predicted highlights per episode for each method. We report the standard deviation in parentheses.
Figure 4: The average number of predicted highlights ($|P|$) for each subset grouped by the number of GT highlights ($|G|$). Shaded area represents $\pm$1 standard deviation around the mean.
In Table 4, we compare the average number of predicted highlights ($|P|$) to ensure a fair evaluation of results. While all non-baseline methods produce a lower average $|P|$ than baseline methods, zero-shot prompting with GPT-4o and Gemini still outperforms other baselines. This suggests that the actual performance gap between GPT-4o and the baselines is larger than what the reported numbers indicate. Furthermore, our method achieves significantly better performance than all other approaches, including zero-shot GPT-4o, while maintaining a similar or lower average $|P|$. This supports our claim that our proposed approach is more effective.
Figure 4 further illustrates how each model adjusts its number of predicted highlights. [^5] Three LLM-based methods with a similar average number of predictions are plotted, along with the frequency-based baseline. We report the same set of methods also in Figure 5. The X-axis represents subsets of test examples, grouped by the number of GT highlights ($|G|$), while the Y-axis shows the average number of predicted highlights ($|P|$) for each subset. Ideally, a model should predict the same number of highlights as the GT (dotted line). However, baseline models predict roughly the same number of highlights regardless of $|G|$. Notably, GPT-4o and Gemini also exhibit little variation in $|P|$ across different $|G|$ values, whereas our approach shows a clear upward trend, indicating a better alignment with the GT distribution.
Performance Across Varying Number of GT Highlights
Figure 5: Precision (left) and recall (right) on test set, grouped by the number of GT highlights. In $|G|=0$ subset, we set both metrics to 1 if the model prediction is an empty set and otherwise 0. Shaded area represents $\pm$1 standard deviation around the mean.
To analyze model performance based on the number of GT highlights per episode, we evaluate precision and recall across subsets of episodes grouped by their $|G|$. Figure 5 shows that the fine-tuned Llama model outperforms all the other methods by a large gap in terms of precision and recall in most subsets.

### v2

@@ Abstract
Podcasts have become daily companions for half a billion users. Given the enormous amount of podcast content available, highlights provide a valuable signal that helps viewers get the gist of an episode and decide if they want to invest in listening to it in its entirety. However, identifying highlights automatically is challenging due to the unstructured and long-form nature of the content. We introduce Rhapsody, a dataset of 13K podcast episodes paired with segment-level highlight scores derived from YouTube’s ‘most replayed’ feature. We frame the podcast highlight detection as a segment-level binary classification task. We explore various baseline approaches, including zero-shot prompting of language models and lightweight fine-tuned language models using segment-level classification heads. Our experimental results indicate that even state-of-the-art language models like GPT-4o and Gemini struggle with this task, while models fine-tuned with in-domain data significantly outperform their zero-shot performance. The fine-tuned model benefits from leveraging both speech signal features and transcripts. These findings highlight the challenges for fine-grained information access in long-form spoken media. We release our codes and dataset at https://github.com/younghanstark/rhapsody.
@@ 1 Introduction
Podcast consumption has grown significantly in recent years, prompting research into content-related tasks such as search (Tian et al., 2022), recommendation (Fan et al., 2023; Nadai et al., 2024), and summarization (Jones et al., 2020; Karlgren et al., 2021). However, identifying the most interesting parts (i.e., highlights) of podcast episodes remains relatively underexplored. This is an important problem, as the predicted highlights could help listeners quickly decide whether an episode is worth their time. Despite its various use cases, progress in this area has been limited due to several challenges: podcasts are often lengthy, and the most interesting parts of the podcast are subjective, making it challenging to gather coherent annotations.
In this paper, we present the first study on automatically locating highlights of podcast episodes. Our approach is inspired by the ‘most replayed’ feature of YouTube: a graph displaying frequently replayed parts of videos (referred to as the replay graph) (YouTube, 2022). This feature, which aggregates the views of thousands of users, opened doors for researchers to study video understanding tasks (Duico et al., 2023; Sul et al., 2023; Kim et al., 2025). Compared to such prior work which mainly consider visual features, we focus on the spoken content.
@@
We first propose the podcast highlight detection task, and derive Rhapsody, [^1] Rhapsody: Replay graph, Highlights, Podcasts, Segment-level, Detection. a large-scale dataset (containing 13K podcast episodes from various domains), visualized in Figure 1. This task is long-context in nature because the average podcast episode in our dataset is about $30$ minutes long and contains an average of 5K words. Each podcast is divided into $100$ equal-length segments, of which a subset (with an average of $5$ segments) is identified as being the most-replayed highlights. We convert the replay graph into the subset of these most-replayed segments. Given input features from the podcast episode, such as episode transcript, episode title, or audio features, a model must identify this subset of most-replayed segments.
To tackle this new task, we propose a suite of baseline systems. We first present a text-based BM25 (Robertson & Zaragoza, 2009) approach, which treats highlight detection as the task of retrieving the segments whose transcript summaries (as summarized by an off-the-shelf LLM) have the highest BM25 score to the episode title. We next present zero-shot prompting results for two state-of-the-art large language models (OpenAI, 2024a; Google Deepmind, 2024). Lastly, we present learned models on this task, which uses hidden representation from a fine-tuned LLM and optionally audio features. This model outperforms GPT-4o’s and Gemini’s zero-shot performance by a large margin. Yet, the task remains very challenging, with a maximum hit rate of $49.0\%$ with our best approach.
@@
  • We formally describe the podcast highlight detection task, explain how supervision signals can be generated from replay graphs, and propose a set of automatic evaluation metrics for this task.
@@
  • We benchmark the performance of various methods, including zero-shot LLM prompting and our fine-tuned language models, across text-only and text-and-audio settings.
We release all our data and code to support future research on podcast understanding. Given the limitations of current models, our benchmark can support rich future work, including methods that combine text and audio signals and methods that focus on the long-context ability of LLMs.
@@ 2.1 Task Formulation
We propose a podcast highlight detection task where a model is provided input features derived from the podcast episode and is asked to predict a set of highlight indices $P=\{p_{k}\}_{k=1}^{N_{P}}$ (where $p_{k}\in\{1,2,\dots,100\}$) which is compared to the GT highlight segment indices $G$. We consider both episode-level input features, such as the episode title, and segment-level input features, such as transcript summaries or audio features for each segment.
@@ 2.2 Data Collection Process
We scrape the list of most popular podcast creators from YouTube, [^2] https://www.youtube.com/podcasts/popularcreators yielding 100 channel IDs. Using YouTube Data API v3, [^3] https://developers.google.com/youtube/v3 we then collect metadata for these channels and their playlists. We discard playlists that (1) are not explicitly set as podcasts by their creators and (2) are not in English. We further filter out playlists that predominantly feature visual content, such as music videos, TV series reviews, or reaction videos. We also discard videos for which the replay graph is unavailable. Then, we limit the maximum duration of the episode to an hour to prevent a single data point in a replay graph corresponding to a long span of the audio. Moreover, we exclude episodes for which automatic transcription has more than 10 segments without any detected words. Details on filtering channels and playlists are in Appendix A.1. We download the audio track, replay graph, and YouTube video metadata for each podcast episode. The metadata includes episode-level features such as the title, creator-written description, video categories, and view statistics.
@@
Figure 2: Comparing two settings for defining GT highlights indices $G$. (Left) Naively defining segments with top-$K$ replay scores as highlights. (Right) Introducing $\tau$, a hyperparameter to impose a minimum threshold (dotted green lines). The light blue lines at the beginning of the graphs represent the graph before bias correction (Appendix A.3).
@@
3 Approaches for Predicting Podcast Highlights
We compare the performance of nine systems: random sampling and frequency-based baselines, a BM25-based method, three zero-shot prompting methods with GPT-4o and Gemini, and three fine-tuned LLMs—one with a segment-level binary classification head and the others additionally incorporating two types of audio features. We first describe feature extraction process from raw audio that will be used by systems.
@@ 3.1 Input Data Preprocessing
To map the podcast transcript to the replay graph, which consists of 100 numbers, we split the transcript into 100 segments. We first run the podcast audio through WhisperX (Bain et al., 2023), a version of Whisper (Radford et al., 2022) that outputs a time-aligned ASR transcription providing word-level alignments of the form $\{(w_{j},s_{j},e_{j})\}_{j=1}^{N_{w}}$, where $w_{j}$ represents the $j$th word, $s_{j}$ its start time, $e_{j}$ its end time, and $N_{w}$ the total number of words. Given this WhisperX output, we segment the transcript into 100 equal-duration intervals based on the episode segments $\{c_{i}\}_{i=1}^{100}$. A word $w_{j}$ is assigned to a segment if both $s_{j}$ and $e_{j}$ fall entirely within that segment’s time span. If a word spans two segments, it is assigned to the segment with the greater overlap. We use whisper-large-v2 for the transcription model and WAV2VEC2_ASR_LARGE_LV60K_960H (Baevski et al., 2020) for the alignment model.
@@
We summarize the transcript segments using gpt-4o-mini-2024-07-18 (OpenAI, 2024b) and use the summary output as input textual features to the model. We prompt LLM to process 10 consecutive segments at a time while providing the previous context to avoid providing an extremely long prompt. We provide more detailed information along with pseudocode in Appendix B.1.
@@
To capture emotional characteristics in podcasts, we explore two types of audio features—dominance-valence-arousal (DVA) embeddings and HuBERT features. The DVA model represents emotions along three dimensions—dominance, valence, and arousal—providing a structured framework for analyzing affective states in speech (Russell & Mehrabian, 1977). To learn richer latent patterns, we leverage the hidden representation from the final transformer layer of an off-the-shelf DVA-based emotion recognition model (Wagner et al., 2023). For HuBERT features, we take the hidden representations from the 10th layer of the model, and average-pool them over time. A more detailed explanation of audio feature extraction can be found in Appendix B.2.
@@ 3.2 Models
As heuristic baselines, we consider random sampling and frequency-based baselines. The random sampling method samples $K$ unique elements from the population of 100 integers $\{1,2,\cdots,100\}$. The frequency-based method precomputes the occurrence count of each index across all GT highlight indices in the training set and selects the top-$K$ most frequent ones. In both methods, $K$ is sampled from the distribution of $|G|$, the number of GT indices, in the training set. All results from these baselines were averaged over 1,000 independent experiments.
@@
As a text-only baseline, we adapt BM25 for our task. We treat the summaries of each transcript segment as the corpus and the episode title as the query. We use 3.0 as the BM25 threshold, which is determined through manual tuning on the validation set.
@@
To test state-of-the-art LLMs on this long-context reasoning task, we consider three zero-shot settings: GPT-4o (text-only), Gemini (text-only), and Gemini (text + audio). For the GPT-4o (text-only) and Gemini (text-only) settings, we prompt gpt-4o-2024-08-06 and gemini-2.0-flash-001 with the title of the episode and transcript summaries for each segment, asking them to identify the segment indices that are most likely to be replayed. The model is instructed to rank answer segments in descending order of replay likelihood for proper computation of AP, and chain-of-thought prompting is applied in the prompt. For the Gemini (text + audio) setup, we also provide the full podcast audio. The text-only and the text + audio prompts can be found in Appendix C.2.
Fine-tuned LLMs with Segment-Level Classification Heads
@@
We fine-tune an LLM on Rhapsody using parameter-efficient fine-tuning and replacing the default next-token prediction head with a custom segment-level binary classification head, shown in Figure 3. We develop three models, a text-only model that only utilizes textual features and two text + audio models that utilize both textual and audio features—DVA or HuBERT embeddings. The input to the text-only model is $100$ transcript summaries, one for each of the $100$ segments, with end-of-segment tokens appended after each segment. We extract one LLM hidden representation for each of the $100$ end-of-segment tokens and apply a binary classification head, consisting of a linear layer and a sigmoid activation, to obtain $100$ probability values, each estimating the likelihood that the corresponding segment is a highlight.
We freeze the base model, add QLoRA (Hu et al., 2022; Dettmers et al., 2023) adapters and optimize these adapters as well as the classification head and linear layer of audio features using binary cross-entropy loss. We use Llama-3.2-1B-Instruct (Meta, 2024) as the base model [^4] We use <|reserved_special_token_0|> for end-of-segment token. and apply AdamW optimization with weight decay. We manually tuned hyperparameters on the validation split and selected the configuration that achieved the highest average performance across all evaluation metrics we report. As detailed in Appendix D, the hyperparameters we tune include the batch size, LoRA hyperparameters, learning rate, and prediction thresholds. The resulting configurations yielded approximately 5.6M, 6.7M, and 6.2M trainable parameters for the text-only, text + DVA, and text + HuBERT settings, respectively.
@@ 4.1 Evaluation Metrics
We aim to measure the similarity of a sequence of predicted indices $P=\{p_{k}\}_{k=1}^{N_{P}}$ based on a sequence of GT indices $G=\{g_{j}\}_{j=1}^{N_{G}}$. Since this setting can be viewed as a document retrieval task where $P$ is a set of retrieved documents and $G$ is a set of relevant documents, we adopt five widely used evaluation metrics for document retrieval — hit rate, precision, recall, F1, and average precision (AP). When computing AP, we ensure $m_{g_{i}}\geq m_{g_{j}}$ is satisfied for all $1\leq i<j\leq N_{G}$.
The number of model predictions heavily influences many evaluation metrics, such as precision and recall. Therefore, we tune the prediction thresholds of the BM25 and fine-tuned Llama models on the validation set, aiming to match the average number of predictions across all methods. We ensure that methods for which we cannot control the number of predictions, such as zero-shot prompting of GPT-4o and Gemini, do not produce excessively many predictions, to enable fair comparison with other models.
@@ 4.2 Results
System Hit Rate Precision Recall F1 AP
@@
Fine-tuned Llama (text-only) 0.452 0.202 0.178 0.163 0.134
Fine-tuned Llama (text + DVA) 0.477 0.216 0.187 0.175 0.140
Fine-tuned Llama (text + HuBERT) 0.490 0.266 0.197 0.197 0.153
@@
We report the main experimental results in Table 2. Our experimental results show that even the best-performing model achieves only 0.266 precision and 0.197 F1 score, highlighting the difficulty of podcast highlight detection and the room for improvement beyond current approaches. Yet, this is a subjective task with derived gold labels, thus the upper-bound performance might not be perfect. We still anticipate a large room for improvement given the current performance level.
As expected, random sampling and frequency-based approaches performed the worst. Amongst the zero-shot prompting approaches, the Gemini model performed the best, while GPT-4o performed similarly to the frequency-based baseline, suggesting that it struggles to infer highlight segments from textual content without explicit training. Finally, our fine-tuned LMs with a segment-level classification head trained on in-domain data, despite having only $\sim 6.2$M trainable parameters, significantly outperform all baselines across all metrics. However, the absolute performance remains lacking, emphasizing the need for further advancements in this task.
@@
Incorporating audio features alongside text leads to consistent performance improvements for the fine-tuned Llama model, suggesting that prosodic and acoustic cues provide valuable information for identifying highlights. Yet, the zero-shot Gemini result with audio does not consistently outperform the text-only Gemini result. This suggests that, while there are useful signals in the spoken content, fine-tuning with in-domain data might be needed to leverage such signals. Providing 100 segmented audio clips may contribute to improved zero-shot Gemini performance compared to providing the full episode as input.
Performance on Different Genres and Duration Groups
Method Ent. News & Poli. Comedy Sci. & Tech. Edu. Sports Ppl. & Blogs
Random Sampling 0.240 0.157 0.238 0.227 0.227 0.257 0.211
Frequency-based 0.235 0.141 0.400 0.153 0.220 0.326 0.259
BM25 (text-only) 0.196 0.217 0.150 0.305 0.252 0.369 0.297
GPT-4o (text-only) 0.286 0.173 0.203 0.275 0.237 0.360 0.264
Gemini (text-only) 0.326 0.219 0.316 0.405 0.326 0.432 0.297
Gemini (text + audio) 0.315 0.185 0.233 0.351 0.319 0.360 0.253
Fine-tuned Llama (text-only) 0.489 0.384 0.714 0.313 0.326 0.586 0.527
Fine-tuned Llama (text + DVA) 0.541 0.403 0.714 0.374 0.341 0.441 0.516
Fine-tuned Llama (text + HuBERT) 0.642 0.427 0.820 0.420 0.444 0.649 0.637
Table 3: Average hit rate across subsets of podcast episodes grouped by genre. When multiple categories are listed in the video metadata, the first category is used as the genre.
Figure 4: Precision (left) and recall (right) on test set, grouped by podcast duration. The shaded area represents $\pm$1 standard deviation around the mean.
In Table 3, we present the average hit rate depending on podcast genres. Fine-tuned Llama model with HuBERT features performs the best, regardless of podcast genre. Notably, fine-tuned Llama models generally outperform other models in the ‘Comedy’ genre by a huge gap. It is also interesting that the fine-tuned Llama model without audio signals performs better than the one with DVA features in the ‘Sports’ genre. Additionally, LM-based methods don’t perform as well in the ‘News & Politics’ genre as the other genres.
In Figure 4, we group podcast episodes by rounding up the durations to the nearest 10 minutes. [^5] Three LLM-based methods with a similar average number of predictions are plotted, along with the frequency-based baseline. We report the same set of methods also in Figure 5 and Figure 6. We do not observe a clear trend; most systems perform similarly regardless of podcast duration.
Performance Across Varying Number of GT Highlights
Figure 5: Precision (left) and recall (right) on test set, grouped by the number of GT highlights. In $|G|=0$ subset, we set both metrics to 1 if the model prediction is an empty set and otherwise 0. The shaded area represents $\pm$1 standard deviation around the mean.
To analyze model performance based on the number of GT highlights per episode, we evaluate precision and recall across subsets of episodes grouped by their $|G|$. Figure 5 shows that the fine-tuned Llama model consistently outperforms all the other methods in terms of precision and recall in most subsets.
@@
FT Llama (text + DVA) 3.96 (3.46)
FT Llama (text + HuBERT) 3.54 (4.09)
Table 4: The average number of GT and predicted highlights per episode for each method. We report the standard deviation in parentheses.
Figure 6: The average number of predicted highlights ($|P|$) for each subset grouped by the number of GT highlights ($|G|$). The shaded area represents $\pm$1 standard deviation around the mean.
In Table 6, we report the average number of predicted highlights ($|P|$) to ensure a fair evaluation of results. The baseline methods, on average, produce a higher number of predictions compared to all non-baseline methods. While zero-shot prompting with GPT-4o and Gemini (text + audio) produces a lower average $|P|$, their performance (as shown in Table 2) is superior to the baseline methods. This indicates that the actual performance gap between them is larger than what the reported numbers indicate. Furthermore, our proposed method achieves significantly better performance than all other approaches, including zero-shot GPT-4o, while maintaining a similar or lower average $|P|$. This supports our claim that our proposed approach is more effective.
Figure 6 further illustrates how each model adjusts its number of predicted highlights. Ideally, a model should predict the same number of highlights as the GT (dotted line). However, baseline models predict roughly the same number of highlights regardless of $|G|$. Notably, GPT-4o and Gemini also exhibit little variation in $|P|$ across different $|G|$ values, whereas our approach shows a clear upward trend, indicating a better alignment with the GT distribution.